OpenStudy (peterg2001):

help will medal Which of the following expressions does not mean the same as the other three? x - 17 x + (-17) x - (-17) - 17 + x @Jaynator495

OpenStudy (peterg2001):

@TheSmartOne @Kit_Kat_Nat_<3

OpenStudy (peterg2001):

@ShadowLegendX

OpenStudy (anonymous):

wat do you think

OpenStudy (anonymous):

its the 3rd one the 17 is positive in that one

OpenStudy (anonymous):

already been answered ahsome stop trying to steal my street cred

OpenStudy (ahsome):

Lets do them one by one. Lets leave the first one the same, and see if the other three are similar: First equation: \(x-17\) The second one is: \(x+(-17)\) When you add a negative number, its the same is minusing the number. Here is an example \(x+(-y)=x-y\) So, \(x+(-17)\) is the same as \(x-17\). That is the same as the first one, so its OK The third one is: \(x-(-17\) When you minus a negative number, its the same is adding the number. Here is an example \(x-(-y)=x+y\) So, \(x-(-17)\) is the same as \(x+17\). This is not the same as the first one, and therefore is the culprit

OpenStudy (ahsome):

That would be the steps, @peterg2001